OpenStudy (lena772):

Sarah wants to buy the car that has a better mileage rating. Cars of Company A travel 20 miles per gallon and cars of Company B travel 14 km per liter. [1 mile = 1.6 km; 1 gallon = 3.79 liters] Part A: Which company's car should Sarah buy? Justify your answer by comparing the mileages after converting to the same units. (5 points) Part B: If 10 gallons of gas is filled in the car which gives better mileage, and you went on a drive and checked on the gas used every 30 miles, describe appropriate scales on the x-axis and y-axis to graph the rate at which the car uses gas. (5 points)

hero (hero):

Hi @Lena772, How far did you get with this one?

OpenStudy (lena772):

Hey. I'm wondering if I should convert km per liter or miles per galon or viceversa

hero (hero):

There are a couple reasons to convert everything to gallons: 1. In the US, most cars and gas stations use the miles per gallon rate. 2. The second question speaks in terms of gallons So it definitely makes sense to use gallons.

hero (hero):

Do you agree?

OpenStudy (lena772):

Yes

OpenStudy (lena772):

14/1.6 = 8.75 miles

hero (hero):

By the way, did you convert km to miles or liters to gallons?

OpenStudy (lena772):

Km to miles

hero (hero):

The reason why I asked is because we were given a rate 14 km/L We have to convert 14km/L to mi/Gal

hero (hero):

Do you get what I mean here?

OpenStudy (lena772):

yes

hero (hero):

I mean if we were only given km, then we could directly convert that to miles, but that's not what we were given.

OpenStudy (lena772):

right

hero (hero):

In order to convert 14 km/L to mi/Gal, we have to do something like this: \[14 \frac{\text{km}}{\text{L}} \times 3.79 \frac{\text{L}}{\text{Gal}} \times\frac{1}{1.6}\frac{\text{m}}{\text{km}}\]

hero (hero):

Then cancel the appropriate units and multiply the numbers to end up with mi/Gal

hero (hero):

\[14 \frac{\cancel{\text{km}}}{\cancel{\text{L}}} \times 3.79 \frac{\cancel{\text{L}}}{\text{Gal}} \times\frac{1}{1.6}\frac{\text{m}}{\cancel{\text{km}}}\]

hero (hero):

\[\frac{(14)(3.79)}{1.6} \frac{\text{mi}}{\text{Gal}}\]

hero (hero):

You may have never seen this kind of conversion before but it is called unit-rate conversion.

hero (hero):

Basically, after conversion you end up with approximately 33 mi/Gal

OpenStudy (lena772):

like molar ratio kind of?

hero (hero):

Whatever helps you understand it. I'm not a chemistry major so not familiar with the term "molar"

hero (hero):

But I believe now we have enough information to answer the first question.

OpenStudy (lena772):

yeA

hero (hero):

So which car should Sarah buy and why?

OpenStudy (lena772):

Car from company B because it has a better gas mileage.

hero (hero):

Yes, company car B is the obvious choice. Now, how will you answer question B?

OpenStudy (lena772):

I suck at graphs I have NO idea

hero (hero):

At the very least, we should be able to figure out units to use for the x and y axes. What units should we use per axis?

OpenStudy (lena772):

gallons on x?

hero (hero):

Remember, the x-axis represents the independent variable. The y-axis represents the dependent variable.

hero (hero):

Also remember that in general the rate will be y/x

hero (hero):

the unit rate

OpenStudy (lena772):

but both of the variables are changing

OpenStudy (lena772):

so how is one independent

hero (hero):

Think of it this way, we need the gas to be able to go anywhere. Without the gallons of gas, no travel is possible.

hero (hero):

Also, we can fill the car with however many gallons we want without going anywhere. Or we can remove the gas from the car without going anywhere.

hero (hero):

However, we can't travel any miles without the gallons of gas.

hero (hero):

So that should be enough for you to understand why gallons is independent and why miles is dependent.

OpenStudy (lena772):

so like i said gallons on x

hero (hero):

Furthermore, the rate is represented by \[\frac{y}{x} = \frac{\text{mi}}{\text{gal}}\]

hero (hero):

So by default, the fraction mi/gal let's us know which axis to put the units.

OpenStudy (lena772):

Right

OpenStudy (lena772):

what about scales

hero (hero):

So far this is what we know graphically: